HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ELECTRONIC SPECTROSCOPY OF VH AND VD: NEW BANDS IN THE RED REGION

\underline{THOMAS D. VARBERG} , NATHAN T. BRANDES, NICK J. IHRKE and KEVIN A. TOVAR, Department of Chemistry, Macalester College, 1600 Grand

Ave., St. Paul, MN, USA

The first gas-phase spectrum of VH, assigned as the D^{5}Π–X^{5}Δ transition at 654 nm, was recently recorded and analyzed in our laboratory.^{1} We have now identified additional bands of VH (at 715 nm) and VD (at 715 and 770 nm), for which rotational analyses are currently being undertaken. At this stage, the 715-nm band of VH has been assigned as arising from the C′^{5}Δ–X^{5}Δ transition. All three bands are red-degraded and display complex rotational structures owing to the molecule’s high spin multiplicity, rapid spin-uncoupling, large upper state Λ-doubling, and presence of strong local perturbations. We are also recording dispersed fluorescence spectra for many of the excitation transitions, which are a valuable aid for assigning the 15 rotational branches of the main sub-bands of these quintet systems. In each case fluorescence is observed to the v = 1 level of the X^{5}Δ state, enabling us to determine the ΔG_{1/2} vibrational intervals of the electronic ground states of VH and VD. The molecules were produced by sputtering vanadium in a hollow cathode source with a dilute mixture of hydrogen or deuterium in argon. The spectra were recorded by laser excitation spectroscopy at Doppler-limited resolution using a continuous-wave ring dye laser. It is hoped that these results will be useful for astronomical searches for vanadium hydride in colder (L-type) stars within our galaxy.

1doi:10.1063/5.0105844, T. D. Varberg, J. Chem. Phys., 157, 074311 (2022).

p-number: p072  Submitted on Sun, 11 Jun 2023 21:41:56 +0200